### v1

# TAIA: Large Language Models are Out-of-Distribution Data Learners

Abstract
Fine-tuning on task-specific question-answer pairs is a predominant method for enhancing the performance of instruction-tuned large language models (LLMs) on downstream tasks. However, in certain specialized domains, such as healthcare or harmless content generation, it is nearly impossible to obtain a large volume of high-quality data that matches the downstream distribution. To improve the performance of LLMs in data-scarce domains with domain-mismatched data, we re-evaluated the Transformer architecture and discovered that not all parameter updates during fine-tuning contribute positively to downstream performance. Our analysis reveals that within the self-attention and feed-forward networks, only the fine-tuned attention parameters are particularly beneficial when the training set’s distribution does not fully align with the test set. Based on this insight, we propose an effective inference-time intervention method: Training All parameters but Inferring with only Attention (TAIA). We empirically validate TAIA using two general instruction-tuning datasets and evaluate it on seven downstream tasks involving math, reasoning, and knowledge understanding across LLMs of different parameter sizes and fine-tuning techniques. Our comprehensive experiments demonstrate that TAIA achieves superior improvements compared to both the fully fine-tuned model and the base model in most scenarios, with significant performance gains. The high tolerance of TAIA to data mismatches makes it resistant to jailbreaking tuning and enhances specialized tasks using general data.
Figure 1: Performance comparison of various fine-tuning methods under three OOD data mixing scenarios. The target domain is medical knowledge, using Chinese subset of MMedBench [52] as the in-domain training dataset. (a) The dataset is mixed with medical OOD data from CMExam [32], maintaining a total dataset size of 20k; (b) The dataset is mixed with general OOD data from CoT-Collection [24], also keeping the total dataset size at 20k; (c) The dataset includes general OOD data from CoT-Collection, while the size of the in-domain training dataset remains at 20k. As the proportion of OOD data increases, the performance of the vanilla fine-tuning declines significantly, whereas TAIA manages to sustain robust performance in the target domain (details in Appendix E.5).
1 Introduction
Large language models (LLMs) have revolutionized Natural Language Processing (NLP), which has been pretrained on a massive textual corpus and encoded massive world knowledge [8, 1]. These models achieve remarkable zero-shot and few-shot performance across a wide range of tasks [6, 2, 45, 65, 66]. The innovation of instruction tuning, also known as supervised fine-tuning (SFT), has further enhanced the instruction-following capabilities of LLMs [46, 11], simplifying human-LLM interactions. Despite the availability of high-quality data for SFT being limited [7, 86], expanding SFT datasets remains a straightforward method to adapt LLMs for specific tasks [14]. Various SFT datasets, such as Alpaca [49] and Natural Instructions [41, 69], have been manually curated or artificially generated to create more generalized instruction-tuned LLMs.
However, real-world applications of LLMs are diverse [6] and complex [34], often making public datasets insufficient. While synthetic data is useful, it is expensive and tends to exhibit a distribution shift biased towards the parent LLM [67]. Consequently, the data distribution that LLMs adapt to during fine-tuning often differs significantly from that required for specific tasks. This discrepancy leads to inferior performance on specialized tasks and knowledge forgetting due to disruptions in the parametric knowledge stored in LLMs [14]. Figure 1 also shows that with more out-of-distribution (OOD) tuning data, the vanilla fine-tune method brings catastrophic forgetting problems, degrading models’ performance on downstream tasks. The scarcity of natural data and the suboptimal quality of synthetic data present substantial challenges to effectively adapting LLMs for specialized tasks. In essence, the dependency on in-domain distribution fine-tuning corpora hampers the broader deployment of LLMs.
To address this, we propose avoiding such data dependency by leveraging the intrinsic properties of fine-tuning and developing an inference-time method that does not rely on high-quality in-distribution data. We first conduct an in-depth investigation of the internal Transformer architecture. We find that during fine-tuning, LLMs enhance their instruction-following ability, primarily controlled by the self-attention module [75]. Conversely, most parameterized knowledge is encoded in the feed-forward network (FFN) module [18, 40] during pretraining [56]. Fine-tuning primarily elicits this pretrained knowledge [71, 46, 59], which remains relatively fixed [86]. This insight prompts us to discard the FFN updates during fine-tuning, as only a small portion positively contributes to downstream performance, while most disrupt the knowledge when fine-tuned on task-mismatched data.
A naive approach is to fine-tune only the attention parameters, but this fails to generalize to OOD data due to insufficient exploration of non-linearity. To ensure sufficient learning of non-linearity, we introduce additional FFN parameters during fine-tuning but retain only the beneficial self-attention updates. This strategy, named Training-All-Inferring-only-Attention (TAIA  [^1] Code is available in https://github.com/pixas/TAIA_LLM), achieves both OOD generalization and sufficient optimization space. The comparisons between the proposed method and the vanilla fine-tuning method are shown in Figure 2
We validate TAIA across seven datasets including math, reasoning, and knowledge understanding, using four LLM families and two fine-tuning techniques. Extensive experiments demonstrate the efficacy of TAIA across various model configurations and its outstanding robustness compared to other baselines. Furthermore, detailed analyses confirm the reproducibility of TAIA in terms of fine-tuning methods and fine-tuning dataset scales. TAIA also maintains the few-shot adaptation ability of base models and withstands multi-level red-teaming attacks. It consistently improves performance in vertical domains like healthcare with increasing OOD data (see Figure 1).
Overall, we conclude our contributions as three-fold:
  1. 1.
    Necessity Analysis: We analyze the necessity of leveraging OOD data for effective downstream fine-tuning, revisiting the roles of self-attention and FFN in the Transformer architecture and formalizing their contributions during fine-tuning.
  2. 2.
    Inference-Time Intervention: We propose a simple yet effective inference-time method that trains all parameters but retains only the self-attention updates. This approach optimizes performance across downstream and closed-book tasks, as validated by extensive experiments.
  3. 3.
    Expanding Model Adaptability: Our approach introduces an innovative method for utilizing OOD data in fine-tuning LLMs, substantially decreasing the dependence on in-domain data. This advancement enhances the adaptability of LLMs, enabling them to perform exceptionally well across a wider range of specialized and real-world tasks.
Figure 2: Comparison between different fine-tuning and inference methods. Parameters colored with green and yellow represent models finetuned with in-domain and out-of-distribution data, respectively. “ID” and “OOD” represents in-distribution and out-of-distribution, respectively.
2 Preliminaries
Self-attention module
Let $\{t_{i}\}_{i=1}^{N}$ represents the inputs to an Transformer-based LLM, and $\{\mathbf{x}_{i}\}_{i=1}^{N}\in\mathbb{R}^{d}$ represent the token representations after the embedding layer of a Transformer-based LLM. For each layer $l$, LLM initially computes the query, key, and value vectors:
$\mathbf{q}_{m}^{l}=f_{q}(\mathbf{x}_{m}^{l},m)\quad\mathbf{k}_{n}^{l}=f_{k}(\mathbf{x}_{n}^{l},n)\quad\mathbf{v}_{n}^{l}=f_{v}(\mathbf{x}_{n}^{l},n)$
where $m,n$ are token indexes in the sequence and $f_{\{q;k;v\}}$ are position embeddings parameterized by RoPE [61]. After that, the attention score is computed between these two position tokens:
$\mathbf{o}_{m}^{l\top}=\mathrm{softmax}\left(\frac{\mathbf{q}_{m}^{l\top}W_{q}^{l}W_{k}^{l\top}\mathbf{K}_{m}^{l\top}}{\sqrt{d_{k}}}\right)\mathbf{V}_{m}^{l}W_{v}^{l}$ (1)
where $W_{q}^{l},W_{k}^{l},W_{v}^{l}\in\mathbb{R}^{d\times d_{k}}$ are learnable weight matrices, $d$ is the model dimension and $d_{k}$ is the inner dimension and $\mathbf{K}_{m}^{l}=[\mathbf{k}_{1}^{l},\cdots,\mathbf{k}_{m}^{l}]^{\top}\in\mathbb{R}^{m\times d},\mathbf{V}_{m}^{l}=[\mathbf{v}_{1}^{l},\cdots,\mathbf{v}_{m}^{l}]^{\top}\in\mathbb{R}^{m\times d}$ and we use the single-head notation for simplicity. Finally, another projection matrix $W_{o}^{l}\in\mathbb{R}^{d\times d_{k}}$ is used to project $\mathbf{o}_{m}^{l}$ back to token space $\mathbf{x}_{m}^{l}=W_{o}^{l}\mathbf{o}_{m}^{l}$. Self-attention with rotary position embedding is more effective for computing contextual mappings [84] in arbitrary sequences, particularly in long contexts [61]. It incorporates an induction-head mechanism that enables the Transformer architecture to predict co-occurring tokens within a given sequence [15, 44] from an in-context perspective. Meanwhile, Wu et al. [75] systematically demonstrate that self-attention significantly enhances its instruction-following capability through fine-tuning. Knowledge tokens that do not appear in the context are stored as global tokens in the FFN memory [5].
Feed-forward network (FFN)
In modern transformer architectures, the SiLU [55] gating linear unit [58] is adopted by various models [65, 66, 3]. It is formulated as:
$\mathbf{x}_{m}^{l}=W_{d}^{l}(\mathrm{SiLU}(W_{g}^{l}\mathbf{x}_{m}^{l})\odot W_{u}^{l}\mathbf{x}_{m}^{l})$ (2)
where $W_{g}^{l},W_{u}^{l}\in\mathbb{R}^{d^{\prime}\times d}$, $W_{d}^{l}\in\mathbb{R}^{d\times d^{\prime}}$ and $d^{\prime}$ is the hidden dimension. $\odot$ is the element-wise multiplication and $\mathrm{SiLU}(x)=x\odot\mathrm{sigmoid}(x)$. The feed-forward network uses inverse-bottleneck parameterization methods, which inherently enlarges the representation space and eases the storage of diverse knowledge from different tasks [18, 40].
Finetuning towards tasks
During fine-tuning in task-related data, natural practices format data as {instruction,(input),output} pairs: $(I,x^{t},y^{t})$ where $t\in\{1,2,\cdots,N\}$ and $N$ is the dataset size. In a causal LLM architecture, the learning objective is to minimize the task distribution with the LLM’s internal distribution, via a negative log-likelihood manner:
$\mathcal{L}_{\boldsymbol{\theta}}=-\frac{1}{N}\sum_{t=1}^{N}\sum_{i=1}^{T}\log p_{\boldsymbol{\theta}}(y_{i}^{t}|y_{<i}^{t},x^{t},I)$ (3)
where $T$ is the output sequence length. This objective prompts $\boldsymbol{\theta}$ to converge when the generated response $\hat{y}^{t}$ matches $y^{t}$, i.e., the internal distribution of $\boldsymbol{\theta}$ aligns with the fine-tuning dataset.
3 TAIA: Training All Parameters but Inferring with Only Attention
3.1 Motivation
Fine-tuning LLMs for downstream tasks typically requires a substantial amount of high-quality conversational data. While a large volume of high-quality synthetic data generated by GPT-4 [45] is publicly available and useful for general domains like Math Word Problems and programming, real-world applications of LLMs are diverse [6] and complex [34]. This diversity renders public datasets insufficient for many scenarios. Synthetic data, although useful, is costly and often exhibits distribution shifts towards the parent LLM [67]. Consequently, the data distribution achieved through fine-tuning often diverges from that required for specific tasks. The scarcity of natural task-specific data and the suboptimal quality of synthetic data present significant challenges for the effective transfer of LLMs to specific tasks.
To address these issues and reduce LLMs’ dependency on specialized data, we aim to enable LLMs to perform proficiently on specific tasks using OOD data. Given that the self-attention and feed-forward network (FFN) modules within LLMs function differently, we re-evaluate their respective roles during supervised fine-tuning. Our study indicates that OOD fine-tuning introduces noisy parametric knowledge into the FFN memory. Therefore, filtering out noisy parameters while retaining beneficial ones is crucial for generalization (§3.2). Through systematic analysis and empirical validation, we demonstrate that the optimal parameter selection strategy is achieved by TAIA, which involves training all parameters but retaining only attention updates (§3.3).
3.2 Parameter Selection for Out-of-Distribution (OOD) SFT
Prior research has demonstrated that LLMs possess a wide range of task knowledge [50, 6, 56] after semi-supervised learning on web data. To enhance their proficiency in specific tasks, domain-related instruction-tuning [85] is utilized to improve the knowledge access process [71, 46, 59]. Moreover, studies [75] have shown that self-attention improves LLMs ability to follow instructions through fine-tuning, which aids in effective knowledge elicitation. However, when trained on OOD data, the optimization objective (Eq. 3) involves significant distribution shifts in certain parameter groups. This can disrupt the pre-trained knowledge stored in the Transformer’s feed-forward network (FFN) [5]. Therefore, a balanced approach is to disregard the parameters that are noisily disrupted, while preserving the parameters that contribute to held-out tasks. This approach is already endorsed by existing research [80]. Since fine-tuning prioritizes effective instruction following over absorbing potentially misleading knowledge, it can be inferred that the knowledge acquired by the FFNs during fine-tuning could be considered somewhat redundant.
3.3 Towards Optimal Parameter Selection Strategy
Based on the above analysis, a subsequent action is to directly fine-tune only the parameters of the self-attention modules and freeze the FFN ones, which we call TOA (Train Only Attention Parameters).
Figure 3: Performance of TOA and TAIA with the layer-wise FFN LoRA. All models are equipped with attention LoRA at each layer and fine-tuned on a corpus mixture with 50% OOD data.
A similar practice to TOA is parameter-efficient fine-tuning (PEFT), as they both only train partial parameters. PEFT has achieved a trade-off between performance and training efficiency due to the superfluity of parameters in LLMs [21, 80]. We anticipate that TOA could yield results comparable to those obtained by training all parameters, as it essentially represents a form of the PEFT method and has been verified in vision tasks [79]. However, as shown in Equation 1, there are few non-linear operations during the attention computation, which inhibits TOA from learning complex representations of the training data. Without sufficient representation exploration, TOA suffers from unlearning of general features via OOD data, despite circumventing catastrophic forgetting problems. We conduct experiments (details in Appendix E.6) to validate the inferiority of TOA in learning general instruction-following ability and show the results in Figure 3. With FFN modules participating in gradient descents, the performance on the downstream task increases with the proper selection of FFN modules, even if the introduction of FFN modules disrupts pretrained memory. However, TOA still lags far behind the base model, indicating its inability to extract non-intervention features from OOD data. To maintain the advantage of the non-linearity of FFN modules on the update of self-attention modules, we adopt another approach: we add all parameters into the optimizer group. In contrast to the vanilla method, we only maintain the updated self-attention part and reuse the pretrained weights for FFN modules after fine-tuning, which we name Training All parameters but Inferring with only Attention (TAIA). TAIA guarantees that self-attention can leverage the gradient descent process to optimize its parameters with redundant FFN fine-tuning parameters. The removal of updated FFN parameters during inference, on the other hand, ensures the integrity of parameterized knowledge stored in original FFN modules and the well-learning of beneficial knowledge from OOD data, as supported in Figure 3.
Table 1: Validation experiments using two training corpus and four seed backbones across seven test sets. “CQA.” refers to CommonsenseQA and “MMB.” refers to the English subset of MMedbench benchmark. “FT Method” denotes the fine-tuning method, which is either LoRA or MoLoRA. Bold indicates the optimal result in each subgroup and underline indicates the suboptimal result. The TAIA setting achieves optimal fine-tuning results in most cases.
Training Dataset Model FT Method Infer Mode Reasoning Knowledge Avg.
MATH BBH CQA. LogiQA SVAMP MMB. MMLU
Base Model Qwen1.5-1.8B 4.28 16.80 58.39 32.41 24.80 33.78 43.62 30.58
Qwen1.5-7B 20.30 30.76 78.30 42.70 54.90 45.09 57.69 47.11
LLaMA2-7B 8.22 26.36 48.40 33.95 44.50 32.21 42.30 33.71
LLaMA3-8B 27.92 29.58 73.71 41.47 83.90 60.33 59.38 53.76
Alpaca- GPT4 Qwen1.5-1.8B LoRA Vanilla 8.02 28.80 60.44 35.02 22.10 34.80 41.67 32.98
TAIA 10.82 30.03 64.29 33.03 29.90 34.64 43.73 35.21
MoLoRA Vanilla 8.44 23.67 60.69 32.72 25.20 34.33 42.58 32.52
TAIA 10.20 27.71 63.47 32.87 31.10 34.33 43.48 34.74
Qwen1.5-7B LoRA Vanilla 17.90 36.09 77.31 37.33 57.10 44.85 54.89 46.50
TAIA 24.98 43.46 77.31 41.78 67.20 46.90 57.29 51.27
MoLoRA Vanilla 16.12 35.05 76.90 38.71 56.80 45.56 55.03 46.31
TAIA 25.04 42.54 77.56 41.94 65.60 46.58 57.15 50.92
LLama2-7B LoRA Vanilla 7.08 33.19 63.96 35.64 43.40 38.02 43.29 37.80
TAIA 8.02 31.47 63.06 34.25 57.08 38.10 41.91 39.13
MoLoRA Vanilla 7.42 32.64 64.95 34.41 45.40 37.23 41.18 37.60
TAIA 8.82 30.93 63.80 34.10 51.80 36.53 41.21 38.17
LLama3-8B LoRA Vanilla 25.26 37.35 75.68 37.63 69.70 57.50 60.84 51.99
TAIA 28.34 31.30 75.92 40.55 85.10 58.68 61.87 54.54
MoLoRA Vanilla 25.38 35.85 77.15 39.63 71.20 57.97 61.78 52.71
TAIA 28.16 29.10 77.48 40.09 84.90 59.07 61.42 54.32
CoT- Collection Qwen1.5-1.8B LoRA Vanilla 7.68 13.90 58.07 21.97 39.00 27.65 25.51 27.68
TAIA 9.64 21.93 67.32 32.57 40.30 34.64 42.39 35.54
MoLoRA Vanilla 7.90 12.99 58.80 22.43 38.20 27.42 23.88 27.37
TAIA 9.08 22.49 67.73 34.56 44.80 36.68 44.13 37.07
Qwen1.5-7B LoRA Vanilla 13.22 24.07 72.65 21.35 53.60 27.49 25.00 33.91
TAIA 20.28 32.54 80.10 41.93 61.90 46.74 56.52 48.57
MoLoRA Vanilla 13.38 22.50 75.59 22.73 52.70 27.57 25.37 34.26
TAIA 19.74 30.96 78.84 41.94 58.81 46.03 57.01 47.62
LLama2-7B LoRA Vanilla 7.98 19.09 56.43 30.57 52.50 38.73 46.99 36.04
TAIA 8.44 26.00 60.77 31.80 58.33 38.33 42.54 38.03
MoLoRA Vanilla 4.54 20.21 61.55 36.26 56.00 37.86 45.09 37.36
TAIA 8.04 30.20 63.49 33.33 55.40 37.55 45.34 39.05
LLama3-8B LoRA Vanilla 16.12 23.24 67.98 27.19 78.60 55.30 60.60 47.00
TAIA 26.28 18.86 71.25 41.17 82.80 58.68 63.16 51.74
MoLoRA Vanilla 17.70 22.24 71.74 28.27 79.30 58.44 60.07 48.25
TAIA 25.46 28.63 73.22 40.40 83.10 60.02 61.82 53.24
4 Experiments
4.1 Backbone LLMs
We select two LLM families, Qwen1.5 [3] and LLaMA [66] and delicately chose control groups to address the following three concerns: (1) Different Model Sizes within the Same LLM Family: We choose Qwen1.5-1.8B and Qwen1.5-7B to test within the same LLM family, how TAIA works for different sizes of LLMs with the same pretraining data; (2) Same Model Size across Different LLM Families: We choose Qwen1.5-7B and LLaMA2-7B to test among different LLM families but the same size, whether TAIA still holds; and (3) Impact of Enlarged Pretraining Data: We choose LLaMA2-7B and LLaMA3-8B to test whether TAIA is applicable when the LLM pretraining data is significantly enlarged. We choose the chat version for all models.
4.2 Experiment Details
We choose two instruction tuning corpus to further demonstrate the high generalization of TAIA under PEFT methods. We choose Alpaca-GPT4-bilingual mixed from Alpaca-GPT4 and Alpaca-GPT4-zh [49]. Apart from this, we also adopt CoT-Collection [24] which is a mixture of various tasks presented in the Chain-of-Thought [72] format. We train 1 epoch for each dataset with the maximum context set to 3072 and the batch size set to 128. We set the learning rate to $42e-$ 4 for all runs and adopt LoRA [21] and Mixture-of-LoRA (MoLoRA)[76, 30] as representative PEFT methods. The LoRA rank is set to 16 and LoRA alpha is set to 32. In MoLoRA, we set the expert count to 4 and activate 1 during inference for all settings. All experiments were conducted on 4 NVIDIA A100 GPUs with ZeRO3 [53] optimization. For the test set, we selected seven widely used datasets: two for evaluating models’ knowledge understanding and five for testing LLMs’ reasoning ability. A detailed description of these test sets can be found in Appendix E.
4.3 Quantitative Analysis
Table 1 presents a comprehensive comparison between various fine-tuning methods (vanilla, LoRA, MoLoRA, and our proposed TAIA) across different training datasets and model backbones. The results clearly demonstrate that TAIA enhances the utilization of training data, consistently outperforming other methods across mentioned benchmarks. For weaker LLMs like Qwen1.5-1.8B and LLaMA2-7B, TAIA significantly amplifies the improvements achieved by standard fine-tuning. For stronger backbones such as Qwen1.5-7B and LLaMA3-8B, standard fine-tuning often degrades performance, but TAIA maintains and even enhances the original capabilities. Notably, TAIA-fine-tuned LLaMA3-8B excels in SVAMP and MMB benchmarks, achieving top scores of 85.10 and 59.07, respectively, indicating its robustness in deep math comprehension and medical reasoning tasks. Furthermore, in the MMLU benchmark, TAIA-fine-tuned models achieve superior average scores, confirming that TAIA not only protects pretrainpretrained knowledge from disturbance but also enables better knowledge utilization for reasoning. These findings underscore the superior efficacy of TAIA in enhancing LLMs’ performance across diverse reasoning and knowledge domains.
Table 2: Comparison with other OOD generalization methods. TAIA is more robust and general than other competitive methods and requires no additional implementation efforts.
Datasets Infer Mode Reasoning Knowledge Avg.
MATH BBH CQA. LogiQA SVAMP MMB. MMLU
Base Model Vanilla 4.28 16.80 58.39 32.41 24.80 33.78 43.62 30.58
Alpaca- GPT4 Vanilla 8.02 28.80 60.44 35.02 22.10 34.80 41.67 32.98
L2 3.68 24.37 57.82 35.33 21.30 35.04 41.30 31.26
EWC 3.56 25.02 60.52 34.10 22.50 34.88 41.07 30.10
Self-Distill 7.34 26.29 53.07 28.57 18.20 35.04 39.87 28.09
LoRACL 8.04 28.80 60.03 34.41 27.40 35.27 42.18 33.73
TAIA 10.82 30.03 64.29 33.03 29.90 34.64 43.73 35.21
CoT- Collection Vanilla 7.68 13.90 58.07 21.97 39.00 27.65 25.51 27.68
L2 0.12 5.66 23.26 22.12 41.50 27.73 23.63 20.64
EWC 0.10 7.71 22.64 22.27 40.40 27.73 23.67 20.65
LoRACL 7.68 14.85 58.07 21.97 41.60 27.65 23.53 27.91
TAIA 9.64 21.93 67.32 32.57 40.30 34.64 42.39 35.54
4.4 Compare with Other OOD Generalization Methods
We mainly choose methods aimed for continual learning (CL) which also attempts to improve models with incoming OOD training data. We select L2, EWC [25], Self-Distill [78] and LoRACL, which is a variant of AdapterCL [39] as the competitors and the detailed settings of the experiment are discussed in Appendix E.7 Table 2 shows that although CL-based methods can leverage OOD data for downstream tasks, they are ineffective in certain evaluation sets (e.g., L2 on MMedBench or LoRACL on CommonsenseQA). It indicates that these methods have specific preferences for downstream tasks and cannot be perfectly applied to any arbitrary application. In contrast, TAIA is not only implementation friendly but also generalizable enough for improving most downstream performances.
4.5 Ablation Study
We test three variants of TAIA, all designed to reduce distribution shifting after fine-tuning on OOD data: TOA, TOF, and TAIF. The latter two, TOF and TAIF, are similar to TOA and TAIA respectively, but with relevant parameters changed from self-attention to FFN. Experiments were conducted on the Qwen1.5-1.8B model using the same setting described in §4.2. Results are shown in Table 5. We observe that both TAIA and TAIF demonstrate better generalization properties compared to the vanilla method, with TAIA achieving the best. This again confirms the crucial role of self-attention in maintaining the generalization ability of LLMs. In contrast, TOF and TOA both suffer from inadequate parameter exploration and even perform worse than the baseline when tuned on OpenMath [64], further supporting the practice of retaining redundant parameters during training.
4.6 Representation Analysis
In §3.3, we infer that TAIA can obtain more general hidden representations compared to the baseline and TOA. We here examine the generalization of TAIA from a perspective of activation similarities. We define the activation similarity of the $i$-th data sample between two models, $\boldsymbol{\theta}_{p}$ and $\boldsymbol{\theta}_{q}$, which are trained separately on two corpora $D_{p}$ and $D_{q}$ with different distributions, as
$\mathrm{Sim}(\mathbf{h}_{p},\mathbf{h}_{q})_{i}=\frac{1}{L}\sum_{l=1}^{L}\frac{\mathbf{h}_{pi}^{l\top}\mathbf{h}_{qi}^{l}}{\|\mathbf{h}_{pi}^{l}\|\cdot\|\mathbf{h}_{qi}^{l}\|}$ (4)
where $\mathbf{h}_{p},\mathbf{h}_{q}$ are the activation hidden states after certain modules, and $L$ is the number of hidden layers. We select C-Eval [22] as the test and Medical-Collection, a 180K subset of CoT-Collection and OpenMath [64] as our training corpus. We follow the same experimental setting as described in §4.2 and present the performance-similarity relations in Figure 5. The results show that a large proportion of activation similarities for TAIA are close to 1, significantly higher than those of other methods. This high activation consistency of TAIA correlates with its superior performance, regardless of the training corpus used. It confirms that by emphasizing instruction-following ability through TAIA, LLMs demonstrate robust generalization performance and effective transferability of training data.
5 Discussion
In this section, we discuss the following research questions (RQ) of the TAIA strategy:
  1. RQ1:
    Does TAIA suit full fine-tuning where the catastrophic forgetting is even more severe?
  2. RQ2:
    We have confirmed that TAIA learns only the beneficial parts of the fine-tuning data. Does this mean that it can survive in red teaming and enhance the model’s helpfulness?
  3. RQ3:
    How proficient can TAIA show if we scale the training corpus?
  4. RQ4:
    Wang et al. [68] finds that supervised fine-tuning hurts LLMs few-shot performance on unseen tasks. Can TAIA restore similar few-shot ability as the base LLM?
  5. RQ5:
    Fine-tuning converges to the downstream distribution, leading to the diminishing rank compared to the base LLM. How does the rank change when TAIA is adopted?
Table 3: The application of TAIA on full fine-tuning technique trained on CoT-Collection. It still surpasses the vanilla fine-tuning method but lags behind the base LLM.
Model Infer Mode MATH BBH CQA. LogiQA SVAMP MMB. MMLU Avg.
Qwen1.5-1.8B Base Model 4.28 16.80 58.39 32.41 27.90 33.78 43.62 31.03
Vanilla 6.88 14.51 59.21 20.28 34.30 27.65 23.36 26.60
TAIA 8.22 15.56 60.61 25.65 39.00 28.20 25.65 28.98
Qwen1.5-7B Base Model 20.30 30.76 78.30 42.70 54.90 45.09 57.69 47.11
Vanilla 9.34 23.85 71.66 21.04 57.90 27.57 24.44 33.69
TAIA 14.60 27.32 72.65 33.79 64.50 40.39 35.90 41.31
Table 4: Comparison of TAIA with vanilla fine-tuning on red-teaming resistance. When jailbreaking LLMs on harmful datasets, TAIA harvests lower attack success rates than vanilla fine-tuning on both harmful and benign datasets, showing its strong generalization in distilling out harmful features.
[…]
LLaMA2-7B-chat 0.00 0.00 0.00 7.66
Vanilla 84.59 93.27 4.04 7.46
TAIA 8.27 30.77 0.38 9.94
Response to RQ1: TAIA is also applicable to the full fine-tuning technique.
Our analysis and empirical study focused on PEFT scenarios, mitigating catastrophic forgetting. To test TAIA in a full fine-tuning context, we maintained the same experiment settings as with LoRA tuning but lowered the learning rate to $55e-$ 5 for stability and used the CoT-Collection as the fine-tuning corpus. Testing on Qwen1.8b and 7b sizes, the results (Table 3) indicate TAIA maintains superior performance in reasoning tasks (SVAMP, MATH, CommonsenseQA). However, due to extensive parameter modifications during full fine-tuning, TAIA experiences significant catastrophic forgetting in knowledge-intensive tasks (MMLU, MMedBench). Despite this, it still outperforms the vanilla inference method, validating its applicability and generalization in full fine-tuning scenarios.
Response to RQ2: TAIA significantly reduces harmfulness and improves helpfulness.
The analysis and experiments above have demonstrated that TAIA enables LLMs to generalize on OOD data, reducing dependency on data quality. To explore if TAIA can handle training data with harmful information while enhancing LLM usefulness without substantially increasing harmfulness, we followed Qi et al. [51] to red-team LLaMA2-7B-chat using three attack levels and evaluated on Advbench [10]. We used 100 of the most harmful samples from the Anthropic red team dataset [16], 10 identity-shifting samples from Qi et al. [51], and benign data from Alpaca-GPT4 [49]. For models tuned on benign data, we also tested helpfulness on AlpacaEval [29]. Results in Table 4 show that TAIA significantly reduces the attack success rate after tuning on three levels of red-teaming data while gaining higher helpfulness from the benign dataset. This demonstrates that careful parameter selection can distill out unsafe parameters and enhance LLM robustness.
Response to RQ3: TAIA succeeds in varying data sizes.
To validate the efficacy of TAIA across different sizes of fine-tuning datasets, we sampled the CoT-Collection dataset to create six fine-tuning corpora of varying sizes: [1K, 10K, 50K, 100K, 200K, 1.8M]. We used the same experimental settings as described in §4. The results, shown in Figure 4(a), indicate that TAIA achieves higher performance more quickly and with less data, demonstrating a more efficient utilization of OOD data. Additionally, unlike vanilla fine-tuning, which experiences significant performance drops when trained on a 1K dataset, TAIA is minimally affected by the distribution gap between its internal distribution and that of the 1,000 samples. This demonstrates the high robustness and generalization capability of TAIA. The full results are detailed in Table 9.
Response to RQ4: TAIA fully restores the few-shot capability of the base LLM and even improves the performance.
Brown et al. [6] demonstrate the generalization of LLMs as they can adapt to new tasks with few-shot in-context learning. As LLMs are incapable of few-shot learning after SFT on a specific dataset [68], we want to verify whether TAIA can maintain the superb few-shot learning ability.We evaluate the few-shot adaptation of TAIA using the same checkpoint trained with CoT-collection and test it in a 100 subset sampled from MATH [20]. Results in Figure 4(b) show that the vanilla fine-tuning method has lost its few-shot learning capability. In contrast, TAIA has regained the ability to learn contextually as the base LLM, achieving performance leap from demonstrations in an approximately linear manner.
Response to RQ5: TAIA increases the representation rank of self-attention.
Dong et al. [13] denote that the residual rank of the representation highly correlates to the final performance of Transformer models. The residual rank of any hidden state $\mathbf{X}\in\mathbb{R}^{n\times d}$ can be obtained by $\|\mathrm{res}(\mathbf{X})\|_{1,\infty}=\|\mathbf{X}-\boldsymbol{\mu}\|_{1,\infty}$, where $\boldsymbol{\mu}\in\mathbb{R}^{d}$ is the averaged representation and $\|\mathbf{X}\|_{1,\infty}=\sqrt{\|\mathbf{X}\|_{1}\|\mathbf{X}\|_{\infty}}$. To quantify this metric human-friendly, we recompute the ratio between the residual rank and the original rank of $\mathbf{X}$ after each attention module as $\|\mathrm{res}(\mathbf{X}_{l})\|_{1,\infty}/\|\mathbf{X}_{l}\|_{1,\infty}$, where $l=0,1,\ldots,L$. The comparisons between the vanilla method, TAIA and TOA trained and evaluated on MATH, are shown in Figure 4(c). We notice a negligible difference between the baseline and TOA, indicating that simply training the self-attention modules does not affect dealing with OOD data. Notably, TAIA significantly increases the residual rank of activations of self-attention modules across all layers, which promises high expressiveness and generality.
Figure 4: (a)Average performance with different sizes of fine-tuning datasets; (b) The few-shot performance on MATH; (c) The layer-wise residual rank of the hidden states on MATH.
6 Conclusion
We revisit the intrinsic properties of LLM fine-tuning and determine that supervised fine-tuning poses minimal requirements for updated FFN parameters. Building on this insight, we introduce TAIA, an inference-time intervention strategy designed to address data scarcity challenges in real-world applications of LLMs. TAIA adheres to the traditional fine-tuning technique but only retains updated self-attention parameters for inference. This approach demonstrates superior generalization ability across various contexts. The high generality of TAIA enables effective training using a mixture of limited in-domain data and extensive OOD data. This method enhances LLM performance on downstream tasks while filtering out undesired information from the fine-tuning set, such as hallucination interference or the decline of few-shot capability.
Appendix A Future Work
TAIA has succeeded in harnessing OOD data for fine-tuning LLMs, reducing substantial reliance on in-domain data. To improve this generalization, there are two main directions to expand this adaptability. Firstly, we hope to find a minimal set of trainable parameters to guarantee sufficient parameter exploration while reducing distribution aliasing. Just as Figure 3 shows, LLMs tuned with $[0,18)$ layers of FFN gain higher performance than vanilla fine-tuning with TAIA method. Coupled with this observation, it is possible to reduce knowledge overlaps between pretrained parameters and downstream ones through a fine-grained selection of tunable parameters. Secondly, an adaptive parameter maintenance strategy instead of the coarse separation of FFN modules can both improve the generalization of LLMs as well as the adaptation of LLMs on knowledge-intensive tasks. We hope our work can provide inspiration on how to improve the parameter utilization of LLMs to adapt to universal distribution datasets.
Appendix B Limitations
Our experiments are conducted based on the assumption that LLMs have gathered certain task knowledge but cannot well utilize it. In tasks like summarization [42] or reading comprehension [54], TAIA still needs to learn the task knowledge except for instruction following. We follow the same experiment setting as §4.2 and use the XSum [42] training set and SQuAD v2.0 [54] training set to fine-tune both 1.8B and 7B sizes of Qwen1.5 models. Table 6 shows that TAIA is inferior to the vanilla fine-tuning method when it is unfamiliar with the downstream knowledge. However, the gap is relatively small (0.7 on SQuAD and 3.21 R-L on XSum on 1.8b scale model), and TAIA reverses such gap in 7B LLM in SQuAD v2.0 (-0.7$\rightarrow$+1.38). For the performance on XSum, it is believed that the model needs to learn specific domain knowledge, such as writing style and word usage preferences, to achieve greater overlap with the reference and thus obtain a higher Rouge score. This shows that TAIA is still applicable to unfamiliar tasks and is significantly suitable for well-pretrained LLMs with sufficient domain knowledge.
Appendix C Broader Impact
TAIA is designed to optimize the fine-tuned LLMs on downstream tasks after tuning on OOD data, by removing the tuned FFN parameters after the normal fine-tuning process. The potential positive implications imply lower difficulty in applying LLM on data-limited domains, including finance or healthcare, thus increasing the accessibility of LLMs on these scenarios. TAIA also makes positive impacts on strengthening the safety and helpfulness of fine-tuned LLMs, and thus brings positive social benefits. Moreover, TAIA does not introduce additional costs and is deployment-friendly. As such, we do not foresee any immediate negative ethical or societal consequences stemming from our work that are different from those that apply to enabling LLMs with OOD generalization capability.
Table 6: Experiments on two tasks whose knowledge is not fully acquired by base LLMs. TAIA lags behind vanilla fine-tuning methods by a small margin for Qwen1.5-1.8B. However, for the base model with sufficient knowledge like Qwen1.5-7B, TAIA surpasses the vanilla fine-tuning methods.
Model Training Data Infer Mode SQuAD v2.0 XSum (R-1/R-2/R-L)
- Base Model 84.00 14.76/3.35/10.05
Vanilla 91.70 14.06/2.05/11.36
SQuAD v2.0 TAIA 91.00 18.88/3.73/12.94
Vanilla 72.11 34.26/13.04/27.48
Qwen1.5-1.8B XSum TAIA 78.48 31.32/10.51/24.27
- Base Model 92.00 21.90/6.17/15.17
Vanilla 93.89 16.75/2.65/13.15
SQuAD v2.0 TAIA 95.27 25.58/7.66/19.79
Vanilla 77.62 42.23/19.99/34.78
Qwen1.5-7B XSum TAIA 81.79 38.51/16.49/31.02
Appendix D Related Work
Supervised Fine-tuning
Supervised Fine-tuning (SFT) is a general methodology to adapt base Large language models (LLM) to downstream tasks and specific domains. By constructing instruction-input-output pairs on target tasks and training base LLMs on such training data with maximum log-likelihood, open-sourced LLMs pretrained on web data can adapt to various domains via zero-shot or few-shot prompting, including medical [74, 9, 30], programming [38, 28, 35], finance [77, 73, 81], and math word problems (MWP) [83, 82, 36]. Due to the large scale of such domain-specific data, fine-tuning the whole large language model is costly; therefore, parameter-efficient fine-tuning (PEFT) is proposed to achieve comparable downstream performances with negligible fine-tuning consumption. Among PEFT methods, Low-rank Adaption (LoRA)[21] and its variants DoRA [33] are the most successful, introducing two trainable low-rank adapters and significantly saving training resources without imposing inference latency
Challenges and Limitations of Fine-tuning
Fine-tuning is a straightforward method for adapting LLMs to various downstream tasks, but it incurs several significant drawbacks, including hallucination, harmfulness, catastrophic forgetting, and safety concerns. Gekhman et al. [17] indicated that fine-tuning instructs the model to produce factually inaccurate responses, as the training process encourages the generation of information not anchored in its pre-existing knowledge base. Furthermore, supervised fine-tuning for specific tasks often results in catastrophic forgetting of the initial alignment [37] and creates trade-offs between helpfulness and harmlessness [4]. Additionally, Kumar et al. [26] highlighted that fine-tuning significantly reduces the resistance of LLMs to jailbreaking, thereby increasing their vulnerability. Qi et al. [51] also demonstrated that even when benign fine-tuning datasets are used, well-aligned LLMs inevitably become more unsafe and harmful, not to mention the issues arising from red-teaming tuning data. In contrast, TAIA can significantly mitigate these drawbacks while still enhancing helpfulness through fine-tuning. By focusing on the intrinsic properties of fine-tuning and developing an inference-time method that leverages only beneficial self-attention updates, TAIA provides a robust solution to the challenges posed by traditional fine-tuning approaches.
Appendix E Experiment Details
E.1 PEFT Settings
For LoRA method, we add a LoRA module for each linear layer except the language model head and embedding layer, resulting in the following target modules: [q_proj, k_proj, v_proj, o_proj, gate_proj, up_proj, down_proj]. For MoLoRA method, we add a vanilla LoRA module for each linear layer in attention modules and an MoLoRA module for each linear layer in FFN modules. This results in the following attention targets: [q_proj, k_proj, v_proj, o_proj] and FFN targets: [gate_proj, up_proj, down_proj], respectively.
E.2 Test Sets Description
We here describe the used seven evaluation sets:
  1. 1.
    MATH [20] is a collection of challenging competition mathematics problems containing 5,000 problems in the test set. Each problem in MATH has a full step-by-step solution which can be used to teach models to generate answer derivations and explanations.
  2. 2.
    BIG-Bench Hard [62] (BBH) is a collection of 23 challenging tasks from BIG-Bench. The 6,511 problems are the tasks for which prior language model evaluations did not outperform the average human-rater.
  3. 3.
    CommonsenseQA [63] is to test models’ ability to answer questions using only the parameterized knowledge instead of the context knowledge. It contains 1,000 problems sourced from ConceptNet [60].
  4. 4.
    LogiQA [31] collects questions about natural language inference (NLI) and requires models to infer the conclusion based on provided premises. It contains 653 problems for both English and Chinese subsets.
  5. 5.
    SVAMP [48] are much simpler datasets compared to MATH, which both test models’ math problem-solving ability. It contains 1,221 problems which are all solvable with one or two simple equations.
  6. 6.
    MMedBench [52] contains test sets written in six languages for testing models’ capability in healthcare. Its English subset contains 1,273 problems and its Chinese subset contains 3,426 problems. We use MMB./MMed EN and MMed ZH indicates the English and Chinese subset, respectively.
  7. 7.
    MMLU [19] is to measure LLM’s multitask accuracy, which contains 14,421 problems. The test covers 57 tasks including elementary mathematics, US history, computer science, law, and more. To attain high accuracy on this test, models must possess extensive world knowledge and problem-solving ability.
E.3 Fine-tuning Sets Description
  1. 1.
    Bilingual Alpaca-GPT4 is a dataset composed of Alpaca-GPT4 and Alpaca-GPT4-Zh [49], which contains 104k sample data in total.
  2. 2.
    COT Collection [24] is a dataset designed to induce Chain-of-Thought [72] capabilities into language models. While proprietary LLMs excel at generating Chain-of-Thoughts based on prompting, smaller LMs do not have this capability. Thus, by fine-tuning to generate Chain-of-Thoughts, it could acquire such abilities.
  3. 3.
    OpenMaths [64] is a math instruction tuning dataset with 1.8M problem-solution pairs generated using a permissively licensed Mixtral-8x7B model [23]. The problems are from GSM8K [12] and MATH [20] training subsets and the solutions are synthetically generated by allowing Mixtral model to use a mix of text reasoning and code blocks executed by Python interpreter.
  4. 4.
    Medical Collection is a collection of bilingual medical multiple choice question answering data with Rational, composed of the Chinese and English subset of MMedBench [52], CMExam [32], and a subset sampled from MedMCQA [47]. It comprises a total of approximately 160k questions, with half in English and half in Chinese.
  5. 5.
    Xsum [43] is a dataset for the evaluation of abstractive single-document summarization systems. The dataset consists of 226,711 news articles accompanied with a one-sentence summary. The articles are collected from BBC articles (2010 to 2017) and cover a wide variety of domains (e.g., News, Politics, Sports, Weather, Business, Technology, Science, Health, Family, Education, Entertainment and Arts).
  6. 6.
    SQuAD v2.0 [54] is a collection of question-answer pairs derived from Wikipedia articles. In SQuAD v2.0, the correct answers to questions can be any sequence of tokens in the given text. SQuAD v2.0 combines the 100,000 questions in SQuAD1.1 with over 50,000 un-answerable questions written adversarially by crowd workers in forms that are similar to the answerable ones.
E.4 Evaluation Details
Our evaluations contain different types of metrics, including Exact Match (EM) and Multiple Choice Accuracy (Acc). For EM metric, we extract the contents followed by The answer is to reduce evaluation biases. For different datasets, we adopt different evaluation prompts after the original problem description:
  1. 1.
    MATH [20], BIG-Bench Hard [62] (BBH), SVAMP [48]: Please format the final answer at the end of the response as: The answer is {answer}.
  2. 2.
    CommonsenseQA [63]: Let’s think step by step. Please format the final answer at the end of the response as: The answer is {answer}.
  3. 3.
    LogiQA [31], MMedBench [52], MMLU [19]: Please answer with option letter directly, do not output other information.
We use greedy decoding to maintain that all results are reproducible.
E.5 Data Mixing Experiments
We constructed synthetic mixed datasets to investigate the impact of different out-of-distribution (OOD) data ratios on the generalization of TAIA. We selected medical knowledge as the target domain and chose MMedBench as the in-domain data and test set. Additionally, we designed three different mixed data scenarios: a) Mixing within the same domain. We selected CMExam [32], which also pertained to the medical domain, as the OOD data and maintained the mixed dataset size at 20k. b) Mixing across different domains. We chose COT, from the general domain, as the OOD data and kept the mixed dataset size at 20k. c) Mixing under a constant in-domain data size. We again selected COT as the OOD data, but this time we maintained the in-domain dataset size at 20k.
For the implementation of fine-tuning, we adopt mixture-of-LoRA (MoLoRA) with four experts in total and activate one of them for each token during inference. The rank $\alpha$ of each LoRA in the miture-of-expert (MOE) is set to 16 and the scale factor $r$ is set to 32. We test the 1.8b and 7b sizes of Qwen1.5, and the results are shown in Fig. 1. It is found that TAIA surpasses the base LLM in all cases, while TOA suffers from performance degradation as the OOD data proportion increases. Additionally, TAIA can utilize OOD data to a greater extent without being adversely affected by the shifted distribution, thereby reducing the model’s dependency on specialized domain datasets. This is particularly significant for tasks with limited data resources.
E.6 Layer-wise FFN Fine-tuning Experiments
Previous work has found that fine-tuning the FFN module can disrupt the knowledge encoded in the model and that fine-tuning the attention module can enhance the model’s ability to follow instructions. However, we have discovered that without the assistance of the FFN, merely fine-tuning the attention module alone does not achieve the expected results. To demonstrate this, we conducted experiments where all attention parameters were fine-tuned, while simultaneously fine-tuning FFN parameters at various positions and quantities to observe their impact on model performance. We choose the fine-tuning corpus mixture with 50 percent of OOD data used in the case (b) of the data mixing experiments in Appendix E.5. We designed seven experimental setups, including $\{[0,6),[0,12),[0,18),[0,24),[6,24),[12,24),[18,24)\}$, where $[0,6)$ indicates that only the FFN modules at layer 1 to 6 (with index 0 to 5) are fine-tuned. We choose Qwen1.5-1.8B as the seed LLM and adopt the LoRA with $\alpha=16$ and $r=32$. As shown in Figure 3, the performance of TAIA achieves the best with the fine-tuned FFN modules located at layers 1 to 18. Besides, with equal parameter quantities, fine-tuning the FFN in the earlier layers aids in the optimization of the attention layers, whereas the later layers do not yield such an improvement. This shows that the self-attention module functions mostly in early to middle layers, which is also consistent with Li et al. [27].
E.7 Comparing with OOD Generalization Methods
Here, we present the implementation details of other OOD generalization methods to ensure the reproducibility of all comparisons. For LoRACL, we compare the PPL value of the input prompt between the base model and the LoRA-tuned model and pick the model with lower PPL as the evaluation model.
For the consolidation method, the fine-tuning loss can be formulated as the sum of the negative log-likelihood (Eq. 3) and the regularization items:
$\mathcal{L}^{reg}_{\theta}=\mathcal{L}_{\theta}+\lambda\sum_{\theta_{i}\in\{\theta\}}\Omega_{i}(\theta_{i}-\theta_{init})^{2}$ (5)
where $\{{\theta}\}$ is the collection of all tunable parameters, $\theta_{init}$ is the parameter of the base model and $\lambda$ is a balance hyperparameter. Considering that we use the LoRA method to fine-tune the model, the updated parameters $\theta_{i}$ can be calculated as the sum of the LoRA parameter $\Delta\theta_{i}$ and the base model: $\theta_{i}=\Delta\theta_{i}+\theta_{init}$. We adopt two types of consolidation methods (L2 and EWC) as the baseline. For L2, we add an L2 constraint to tunable parameters:
$\mathcal{L}^{L2}_{\theta}=\mathcal{L}_{\theta}+\lambda\sum_{\theta_{i}\in\{\theta\}}\Delta\theta_{i}^{2}$ (6)
For EWC, we use the Fisher information matrix to measure the distance between the parameters:
$\mathcal{L}^{EWC}_{\theta}=\mathcal{L}_{\theta}+\lambda\sum_{\theta_{i}\in\{\theta\}}F_{i}\Delta\theta_{i}^{2}$ (7)
where $F_{i}$ is estimated by calculating the batch average of the squared gradients of the parameters.
For Self-Distill, we first use the official prompt of Yang et al. [78] to generate the distilled Alpaca-GPT4 dataset using Qwen1.5-1.8B and fine-tune it using the distilled dataset with the same experiment setting with vanilla fine-tuning.
Table 7: Comparison of TAIA with vanilla fine-tuning on hallucination resistance. When fine-tuned on datasets with different quality levels, TAIA harvests lower performance drops than vanilla fine-tuning, showing its strong generalization in distilling out hallucinated features.
Infer Mode Qwen 1.8b Qwen 7b LLaMA2 7b LLaMA3 8b
Base Model 4.39 24.11 7.66 26.31
ShareGPT-52K
Vanilla 1.22 1.67 1.99 3.52
TAIA 4.37 9.68 4.87 7.64
Alpaca-GPT4
Vanilla 5.09 10.42 7.46 16.34
TAIA 4.98 11.46 9.94 18.33
Appendix F Supplementary Experiments
F.1 More Results of TAIA on Helpfulness
We here discuss the effect of data quality on TAIA. We select two fine-tuning datasets, ShareGPT-52K [57] generated from gpt-3.5-turbo and Alpaca-GPT4 [49] generated from GPT4 [45]. ShareGPT-52K collects data from both GPT API and websites, so it contains noisy samples or hallucinations, including misspelled words or repeated sentences. On the other hand, Alpaca-GPT4 collects samples generated through Self-Instruct [70], in which the automatic generation process guarantees fewer noisy contents. We aim to verify whether TAIA still benefits from low-quality data. We choose Alpaca-Eval [29] as the evaluation set which uses GPT4 as the evaluator. We use the same hyperparameter settings as §4.2 to train each model family on these two datasets and display the results in Table 7. We observe that the noisy contents bring a significant decrease in helpfulness across all LLM families with the vanilla fine-tuning technique. In contrast, TAIA reduces such performance drops, especially when tuned in the ShareGPT-52K dataset, and even harvests higher helpfulness than base LLMs when they are relatively weak (e.g., Qwen 1.8B and LLaMA2 7B).
F.2 Different LoRA Ranks Experiments
In this section, we explore the impact of varying the LoRA ranks in the attention and FFN modules on the TAIA. We fine-tuned the Qwen1.5-1.8B model on the medical collection dataset and tested its performance on the general knowledge benchmark under different conditions. We use $ar$ and $fr$ to represent the rank of attention and FFN modules, respectively. As shown in Table 8, when the $ar$ is less than or equal to $fr$, the TAIA achieves the best performance among the three cases and significantly surpasses the Vanilla. However, when the $ar$ is larger than $fr$, the performance of the TAIA lags behind the TAIF. This indicates that with more parameter alteration, self-attention will function more as FFN to encode knowledge, which brings significant knowledge overlap with pretrained models and finally results in worse knowledge understanding performances.
F.3 Full Results on Dataset Scales
We present the full experiment results discussed in RQ3 of §5 in Table 9.
Table 8: Experiments of the different ranks of Attention/FFN LoRA. The ranks of attention and FFN module are noted as ‘ar’ and ‘fr’, respectively. For example, the case ‘ar4_fr64’ indicates the attention rank is 4, and the FFN rank is 64. The results show that TAIF will have better performance than TAIA only when the attention rank is much greater than the FFN rank.
Knowledge
Training Data Model Train/Infer CMMLU MMLU C-Eval Avg.
- Qwen1.5-1.8B -/- 52.68 43.62 55.57 50.62
Vanilla 39.60 27.47 37.74 34.94
TAIA 54.58 44.47 55.57 51.54
LoRA ar4_fr4 TAIF 47.06 42.05 45.62 44.91
Vanilla 45.22 27.72 45.32 39.42
TAIA 54.20 43.80 56.32 51.44
LoRA ar4_fr64 TAIF 45.76 29.70 46.14 40.53
Vanilla 45.05 28.79 42.79 38.88
TAIA 51.24 40.26 48.74 46.75
Medical Collection LoRA ar64_fr4 TAIF 54.25 44.33 55.65 51.41
Table 9: Full results of the ablation experiment on fine-tuning data size. We choose six data scales [1k, 10k, 50k, 100k, 200k, 1.8M] to validate TAIA’s effectiveness.
Data Size Infer Mode Reasoning Knowledge
MATH BBH CQA. LogiQA SVAMP MMedB. MMLU Avg.
- Base Model 8.22 26.36 48.40 33.95 44.5 32.21 42.30 33.71
Vanilla 6.70 18.26 56.43 29.19 35.50 26.39 35.86 29.76
1k TAIA 6.50 21.86 60.77 31.80 47.5 27.89 40.00 33.76
Vanilla 7.74 18.06 51.68 34.56 52.80 37.47 44.7 35.29
10k TAIA 8.34 31.64 59.79 33.18 49.10 39.98 44.67 38.10
Vanilla 7.46 19.35 55.86 30.57 52.10 37.71 45.13 35.45
50k TAIA 8.54 32.42 61.67 32.87 49.10 39.67 45.16 38.49
Vanilla 7.08 20.14 59.62 33.64 53.70 38.81 44.93 36.85
100k TAIA 8.02 30.90 63.72 30.72 47.80 37.71 45.57 37.78
Vanilla 7.98 19.09 56.43 30.57 52.50 38.73 46.99 36.04
200k TAIA 8.44 26.00 60.77 31.80 58.33 38.33 42.54 38.03
Vanilla 7.50 15.36 75.68 34.41 65.30 38.10 44.17 40.07
1.8M TAIA 8.08 30.23 66.91 33.03 58.10 39.59 47.78 40.53
Appendix G Case Study
Here we present three cases of TAIA and vanilla method evaluated on Advbench in Figure 67 and 8, which are tuned on explicitly harmful contents, identity shifting contents and benign contents, respectively. TAIA produces harmless contents against severe jailbreaking, albeit undertaking red teaming tuning.
Figure 6: Comparison between TAIA and vanilla fine-tuning method on Advbench with explicitly harmful tuning contents. We choose LLaMA2-7B-chat as the base LLM. We highlight the harmful contents generated by the vanilla method and denote the rejective contents produced by TAIA.
Figure 7: Comparison between TAIA and vanilla fine-tuning method on Advbench with identity shifting tuning contents. We choose LLaMA2-7B-chat as the base LLM. We highlight the harmful contents generated by the vanilla method and denote the rejective contents produced by TAIA.
Figure 8: Comparison between TAIA and vanilla fine-tuning method on Advbench with benign tuning contents (Alpaca-GPT4). We choose LLaMA2-7B-chat as the base LLM. We highlight the harmful contents generated by the vanilla method and denote the rejective contents produced by TAIA.